OpenStudy (anonymous):

Can someone please help me with this question. Thank you so much in advance! :)

OpenStudy (zzr0ck3r):

13

OpenStudy (uri):

No -13

OpenStudy (anonymous):

I didn't post the question yet....

OpenStudy (ja1):

You both are incorrect.. it is indeed 42..

OpenStudy (uri):

@nslesin We post question in the question box..

OpenStudy (zzr0ck3r):

m = 13 a = 1 t = 20 h = 8 m+a+t+h = 13+1+20+8 = 42 @JA1

OpenStudy (zzr0ck3r):

@nslesin what is the question?

OpenStudy (zzr0ck3r):

sorry no idea what that is

OpenStudy (anonymous):

argh

OpenStudy (ja1):

I told you @zzr0ck3r